# Is Our Chatbot Telling Lies? Assessing Correctness of an LLM-based Dutch Support Chatbot

### v1

@@ Abstract
This study is the first to define the correctness of a response based on how the support team at AFAS makes decisions. It leverages literature on natural language generation and automated answer grading systems to automate the decision-making of the customer support team. We investigated questions requiring a binary response (e.g., Would it be possible to adjust tax rates manually?) or instructions (e.g., How would I adjust tax rate manually?) to test how close our automated approach reaches support rating. Our approach can identify wrong messages in 55% of the cases. This work shows the viability of automatically assessing when our chatbot tell lies.
Index Terms:
Trustworthy AI, Validation of AI-based System, Correctness, Chatbot, Large Language Models.
I Introduction
Companies value their customers [2] and strive to create a great customer experience[3]. Customers, in turn, assess a company on its core business and customer service, which influences their trust, loyalty, and satisfaction [4]. Today, the most popular way to assist customers online is via chatbots and live chats [5, 6, 2]. Real-time communication in live chats means quick answers to questions [6, 2], which helps build loyalty[5] and encourages customers to return when they need assistance[6].
With recent advancements in Large Language Models (LLM) that enable natural and chat-like communication [9, 10, 12], AFAS sees an opportunity to provide live support. The first part of Figure 1 represents the current situation. When a customer raises an issue, an employee forwards the question to the chatbot along with relevant documents and instructions, also called a system prompt. When the LLMs generate an answer based on the information provided (details in II), the support employee checks the answer and forwards it to the customer if correct.
@@ I Introduction
Figure 1: shows the current and desired flow for handling customer queries. In the current workflow, the support team is an intermediate for providing context* comprising of relevant documents and instructions for the large language model and later assessing the response (see part blue). Using green parts, we envision replacing human feedback with automated ratings.
Our first challenge in building an automated solution is understanding what makes a response right. To solicit an answer to this question, the first author shadowed a support staff for a day to observe and interactively understand what makes a response right. Combined with the literature search and analysis of reasons for rejecting chatbot responses, this showed that the first step to the ‘right’ answer is correctness, characterized in terms of relatedness, completeness, and truthfulness. With much research focusing on relatedness [18, 15], this study focuses on truthfulness. To assess our approach, in the first round we gathered a data of 79 posts which we used for training. At a later point, we collected data from 154 posts for testing. The limited data size characterize our analysis.
Since the training data is scarce, it is not possible to train a model using reference answers as has been widely seen in literature [13, 14, 15, 16, 17]. As an alternative, we model how the support team makes decisions and derive heuristics. To measure these heuristics, we take inspiration from Natural Language Generation [15, 14, 19] and Automated Answer Grading literature for metrics [17, 16, 45, 71, 46]. In the process, we note that the choice of heuristics varies with the type of question asked. For example, heuristics for assessing the correctness of a yes/no answer are different from the heuristics for a question that solicits instructions.
Our resulting model assesses the correctness of yes/no questions and questions requiring instructions to show that our model can detect a very inaccurate response with 55% accuracy. Notably, the overall accuracy is better for translated text in English than in Dutch. Further, we observed a 0.3 correlation of our score with human evaluation for Dutch text and 0.37 for the translated text in English, both of which is higher than the 0.13 reported by Mehri et al. [20] in their study on generic conversations.
@@
  • a working definition of correctness to evaluate chatbot responses
  • derive heuristics and customized metrics to assess correctness
  • shows relevance of translating regional language text to English for higher accuracy, and
  • lists the importance of question type for the choice of heuristics for assessing correctness.
II Background
@@
AFAS is a software company based in the Netherlands specializing in automating business processes. More than 3 million users have utilized their software, leading to almost 112,000 support inquiries in 2023.111See annual report: https://jaarverslag.afas.nl/2023 Addressing these support queries demands a lot of time and commitment from the 70-member support team. Since the support team receives nearly 1 query every minute about their software, saving time on even a subset of the queries will be helpful.
The chatbot of AFAS uses Retrieval Augmented Generation (or RAG) to improve the performance of the language model. RAG comprises four parts[21]:
@@
  • Indexing — Creates an index of all documents containing relevant information for a user. This step is done before the chatbot serves any answers. At the time of writing, AFAS indexes help documents.
@@
  • Search — Relevant documents are retrieved based on their similarity to the user message using embedding similarity [22] and keyword-matching techniques [23, 24]. Figure 1 shows this step as Find Relevant Documents.
@@
  • Prompting — This step combines the user message, relevant documents, and system prompt as a single message. The system prompt includes instructions and basic information, such as ’be friendly’ and contact information. Since the chatbot relies on an LLM that is not fine-tuned on company data but is a generally trained model, the LLM requires necessary information to provide relevant responses. Relevant documents provide this information to the LLM. This is shown as Question + Context from Chatbot to LLM in Figure 1.
@@
  • Inference — The question + context is used to prompt an LLM, and the generated response is shown to the support team. This is depicted as the answer from the LLM to the chatbot in Figure 1.
@@
We envision the chatbot to handle all kinds of questions, considering the unique jargon of the company/industry and the fact that it mainly serves Dutch users. Here, reference answers may not help since they are sparse and do not ensure a right response to unseen use cases. Given these constraints, there is a need for a generic definition of what makes a right answer and how to measure it.
III What makes a right answer?
@@
There are two ways to assess whether a chatbot gives right answers: turn- and dialogue-level metrics. Turn-level metrics rate a single message-answer pair[15, 25, 26, 14, 27, 28, 29, 30, 19, 20]. In contrast, dialogue-level metrics rate the full dialogue, including all message-answer pairs [31, 32, 33, 34, 35]. Our objective is to assess the correctness of each answer and, hence, turn-level.
@@
To the best of our knowledge, no prior work defines correctness or not clear enough [20] for measurement and validation. Therefore, our first objective was to define correctness. To define correctness, we followed a two-pronged approach. First, we looked at 500 chatbot responses for which the support team provided a decision: accept or reject and a short justification for rejection. Further, the first author shadowed [36] an experienced support employee throughout one workday and interactively discussed the decision-making. The referenced employee has 12 years of experience in various product support areas of the software, has been using chatbot since its launch, and has necessary training (e.g., a course on using ChatGPT). During the discussions, the employee discussed special cases where the chatbot did not meet expectations and what modifications were required to ensure the answer was correct before presenting it to the customer.
@@
Based on shadowing, discussion, and analysis of rejection reports, the three most common mistakes and, hence, requirements for correctness stood out. They are (as defined in the Oxford Dictionary [37]):
@@
”The quality of only saying what is true”,
@@
”A close connection with the subject you are discussing or the situation you are in”,
@@
”The fact of including all the parts, etc. that are necessary; the fact of being whole”.
A response is considered correct if it contains only true information (truthfulness), is related to the situation and question (relatedness), and comprises all relevant information and solutions (completeness). Later, we designed a plugin soliciting support team responses to understand which requirements the chatbots fall short of, as visualized in Figure 1.
Returning to the literature, we observed that relatedness is well-researched[15, 25, 26]. However, truthfulness and completeness are not. Of the remaining two, we study truthfulness since if an answer is not true, completeness would not matter. For an incomplete answer, the customer can ask follow-up questions, but if untrue information is presented to the customer, it can cause harm to the customer and the company. In the future, a combination of the above three dimensions can be used to measure correctness. The rest of the paper measures truthfulness and assesses it with respect to the manually rated ground truth.
IV Methodology
Our methodology has three parts. First, we determine how the support team assesses truthfulness. Second, we translate how the support team assesses truthfulness into metrics. Finally, we determine how well our approach works. Further, since most research and tools used in similar studies are in English, we explore as a side experiment whether the translated text in English performs better than the Dutch text.
@@ IV Methodology
IV-A Data Gathering
@@
Currently, the developed chatbot of AFAS is utilized by the support team as an assistant. They use the bot to answer questions they have themselves or to answer the questions of a customer. Based on their expert knowledge, they rate the answer of the chatbot for truthfulness. They are encouraged to rate messages, but may choose which messages to rate by themselves. Consequently only a few messages are rated each day, this challenge is posed as we work with a real-case company scenario. The rating is along a Likert scale, which ranges from (1) very untrue to (5) very true. Likert scales are commonly used in similar research [39]. While the size of the scale is a topic of debate, a 5-point Likert scale is most commonly used [40, 41, 42]. During the research, the team continued to use the chatbot, which meant that new feedback was continually being received.
@@
During the study, the plugin to rate truthfulness is implemented. A few weeks after the rating option is introduced, the data is extracted from the system to form an analysis set, consisting of 79 samples. A few weeks later, a test set is extracted, containing 154 message-answer pairs. For each rated message-answer pair, the context (relevant documents + system prompt) is gathered as well. Finally, to test whether English text performs better than Dutch, we translate the data using Google Translator [^2] https://pypi.org/project/deep-translator/. Since our raters frequently rate an answer as either fully true or fully untrue, the dataset becomes imbalanced, see Table I. Consequently, we focus less on overall accuracy and more on the accuracy of detecting 1-star and 5-star rated messages. These messages have a larger number and are of greater interest to the company. Since fully untrue answers must not be sent to users, and fully true answers can be sent without human intervention.
@@
Train set (Test set) 22 (32) 2 (5) 8 (19) 6 (16) 41 (82)
TABLE I: Message count for 1-5 star ratings, with test set values in parentheses.
IV-B Constructing the Tree
@@
To ensure that our metric correlates with human ratings, it is essential to understand how a human determines their rating. In order to represent the thought process of human annotators, we sought a model that could be easily visualized and understood. Furthermore, we sought a method to systematically organize the analysis and maintain a record of the observed characteristics. We opted for a decision tree to depict the mental process, with each node symbolizing the subconscious decisions made by the annotator. We use a manual created tree, as our goal is to mimic the human workflow. If we use automated decision tree builders like Random Forests[43] or C4.5[44], they would create their own reflection and would not reflect the human workflow.
@@
In relevant literature, automated decision-makers utilize features that range from simple and syntactic to complex and semantic ones. Simple features might verify the presence of keywords [16, 45, 19], while more complex features involve evaluating the relevance [18, 15] and meaning of a sentence [46, 47, 48]. The decision tree is build up based on this hierarchy, starting with syntactic checks and progressing to nuanced semantic evaluations if the syntactic checks hold. With this approach, the construction and human evaluation process becomes efficient, as it ensures that complex evaluations are not always necessary to be carried out.
@@
The tree is constructed by the first author through iterative testing and modification using message-answer pairs. For the initial message-answer pair, the author assessed what makes the answer untrue and added a node describing the mistake. Then, they evaluated another message-answer pair to check if the decision tree properly rejects the answer. If not, a node is added or changed. This procedure is executed for all message-answer pairs and consistently cross-checked with the pairs that have already been evaluated. As the decision tree is constructed by a single researcher, of course the representation differs if another human manually creates it. However, the created decision tree is confirmed by multiple employees of the company that it correctly reflects the annotation process. The final version of the decision tree can then be used to determine whether an answer is truthful. Although the decision tree does not make perfect decisions, it effectively visualizes the mental process of an annotator. A part of the tree is shown in Figure 2, and the full tree is included in the replication package [1].
@@
Figure 2: A snippet of the decision tree, indicating whether an answer would be true or not.
@@
The first author has no knowledge about the product during the construction of the tree and has relied on the evaluation in the same context as our metric and the LLM. This way, the knowledge of the author is comparable to the knowledge of the LLM and the metric. It is crucial, as the metric, like the author, should be able to calculate a score without a reference answer and, therefore, without relying on external knowledge from such a reference. This is possible since the LLM is not trained, as far as we know on the company documentation. All info in the answer about the product should be contained in the context; otherwise, the LLM will just come up with info on its own. If not in the context, the LLM is assumed to be hallucinating [19].
@@
IV-C Deriving Heuristics from the Decision Tree

### v2

# mode = titleIs Our Chatbot Telling Lies?

@@ Abstract
This study is the first to define the correctness of a response based on how the support team at AFAS makes decisions. It leverages literature on natural language generation and automated answer grading systems to automate the decision-making of the customer support team. We investigated questions requiring a binary response (e.g., Would it be possible to adjust tax rates manually?) or instructions (e.g., How would I adjust tax rate manually?) to test how close our automated approach reaches support rating. Our approach can identify wrong messages in 55% of the cases. This work demonstrates the potential for automatically assessing when our chatbot may provide incorrect or misleading answers. Specifically, we contribute (1) a definition and metrics for assessing correctness, and (2) suggestions to improve correctness with respect to regional language and question type.
keywords:
[…]
1 Introduction
Companies value their customers (Chattaraman et al., 2012) and strive to create a great customer experience (Becker and Jaakkola, 2020). Customers, in turn, assess a company on its core business and customer service, which influences their trust, loyalty, and satisfaction (Parasuraman, 1998). Today, the most popular way to assist customers online is via chatbots and live chats (Shereen, 2024; Turel and Connely, 2013; Chattaraman et al., 2012). Real-time communication in live chats means quick answers to questions (Turel and Connely, 2013; Chattaraman et al., 2012), which helps build loyalty (Shereen, 2024) and encourages customers to return when they need assistance (Turel and Connely, 2013).
With recent advancements in Large Language Models (LLM) that enable natural and chat-like communication (Bhattacharya et al., 2024; Hagos et al., 2024; Santosh et al., 2024), AFAS sees an opportunity to provide live support. The upper half of Figure 1 represents the current situation. When a customer raises an issue, an employee forwards the question to the chatbot along with relevant documents and instructions, also called a system prompt. When the LLMs generates an answer based on the information provided (details in 2), the support employee checks the answer and forwards it to the customer if correct.
@@ 1 Introduction
Figure 1: Shows the current and desired flow for handling customer queries. In the current workflow, the support team is an intermediate for providing context* comprising of relevant documents and instructions for the large language model and later assessing the response (see ’Chatbot’ and ’Large Language Model’). Using parts ’Google Translator’ and ’Automated Rating’, we envision replacing human feedback with automated ratings
Our first challenge in building an automated solution is understanding what makes a response right. To solicit an answer to this question, the first author shadowed a support staff for a day to observe and interactively understand what makes a response right. Combined with the literature search and analysis of reasons for rejecting chatbot responses, this showed that the first step to the ‘right’ answer is correctness, characterized in terms of relatedness, completeness, and truthfulness. With much research focusing on relatedness (Merdivan et al., 2020; Zhang et al., 2020), this study focuses on truthfulness. To assess our approach, in the first round we gathered a data of 79 posts which we used for training. At a later point, we collected data from 154 posts for testing. The limited data size characterize our study.
Since the training data is scarce, it is not possible to train a model using reference answers as has been widely seen in literature (Papineni et al., 2002; Banerjee et al., 2023; Zhang et al., 2020; Kumar et al., 2019; Roy et al., 2016). As an alternative, we model how the support team makes decisions and derive heuristics. To measure these heuristics, we take inspiration from Natural Language Generation (Zhang et al., 2020; Banerjee et al., 2023; Roychowdhury et al., 2023) and Automated Answer Grading literature for metrics (Roy et al., 2016; Kumar et al., 2019; Jamil and Hameed, 2023; Lakshmi and Simha, 2022; Vij et al., 2019). In the process, we note that the choice of heuristics varies with the type of question asked. For example, heuristics for assessing the correctness of a yes/no answer are different from the heuristics for a question that solicits instructions.
Our resulting model assesses the correctness of yes/no questions and questions requiring instructions to show that our model can detect a very inaccurate response with 55% accuracy. Notably, the overall accuracy is better for translated text in English than in Dutch. Further, we observed a 0.3 correlation of our score with human evaluation for Dutch text and 0.37 for the translated text in English, both of which is higher than the 0.13 reported by Mehri and Eskenazi (2020) in their study on generic conversations.
@@
  • a working definition and metrics to assess the correctness of the responses generated by LLMs
  • suggestions to improve the correctness for regional language and type of question
The proposed definition provides a structured foundation for feedback. We decomposed the definition of correctness into smaller, well-defined definitions. Making chatbot evaluation more concrete and manageable. Our methodology further demonstrates how this definition can be utilized in an empirical evaluation. Moreover, the proposed metric and features may be directly applied or serve as foundation for developing customized evaluation metrics for LLMs. Through our contributions, we illustrate how software companies developing chatbots can implement guardrails for their systems. These metrics will help the system to prevent the bot from telling lies and thus improving the quality of the chatbot. Finally, our recommendations for improving correctness in regional languages can serve to further advance the evaluation of chatbots.
2 Industrial Setting
@@
AFAS is a software company that specializes in automating business processes through their ERP system. Its headquarter resides in The Netherlands, but AFAS has offices in Belgium and on the Caribbean too. The software product is provided as a service to thousands of organizations. In 2023 more than 3 million users utilized their software, leading to almost 112,000 support inquiries. [^1] See annual report: https://jaarverslag.afas.nl/2023 These support queries are handled by a support team of 70 people who dedicate their time to this task. Since the support team receives nearly one query every minute about their software, saving time on even a subset of the queries will be helpful.
To grow its user base without proportionally growing the support team, AFAS started to develop a chatbot that automates the answering of support queries. This chatbot is developed by a team consisting of four developers. They use internal developed frameworks for both back-end processing and front-end rendering.
Using LLMs and grounding the prompts with relevant documentation, the AFAS development team hopes to unburden the support team by not only providing them with a possible answer, but also showing them which documentation is relevant. The AFAS chatbot uses Retrieval Augmented Generation (or RAG) to improve the performance of the language model. RAG comprises four parts (Yu et al., 2024):
@@
  • Indexing | Creates an index of all documents containing relevant information for a user. This step is done before the chatbot serves any answers. At the time of writing, AFAS indexes help documents.
@@
  • Search | Relevant documents are retrieved based on their similarity to the user message using embedding similarity (Devlin et al., 2019) and keyword-matching techniques (Ramons, 2003; Robertson and Zaragoza, 2009). Figure 1 shows this step as Find Relevant Documents.
@@
  • Prompting | This step combines the user message, relevant documents, and system prompt as a single message. The system prompt includes instructions and basic information, such as ’be friendly’ and contact information. Since the chatbot relies on an LLM that is not fine-tuned on company data but is a generally trained model, the LLM requires necessary information to provide relevant responses. Relevant documents provide this information to the LLM. This is shown as Question + Context from Chatbot to LLM in Figure 1.
@@
  • Inference | The question + context is used to prompt an LLM, and the generated response is shown to the support team. This is depicted as the answer from the LLM to the chatbot in Figure 1.
@@
We envision the chatbot to handle all kinds of questions, considering the unique jargon of the company/industry and the fact that it mainly serves Dutch users. Here, reference answers may not help since they are sparse and do not ensure a right response to unseen use cases. Unseen questions are expected, as we consider a user-driven chatbot (Følstad et al., 2019). Users can ask a wide variety of questions, which means the metric must be prepared to handle unseen questions. Given these constraints, there is a need for a generic definition of what makes a right answer and how to measure it.
3 What makes a right answer?
@@
There are two ways to assess whether a chatbot gives right answers: turn- and dialogue-level metrics. Turn-level metrics rate a single message-answer pair (Zhang et al., 2020; Singh et al., 2021; Yan et al., 2016; Banerjee et al., 2023; Das and Verma, 2020; Phy et al., 2020; Tao et al., 2018; Gupta et al., 2022; Roychowdhury et al., 2023; Mehri and Eskenazi, 2020). In contrast, dialogue-level metrics rate the full dialogue, including all message-answer pairs (Yeh et al., 2021; Huang et al., 2020; Pang et al., 2020; Deriu et al., 2020; Lowe et al., 2017). Our objective is to assess the correctness of each answer and, hence, turn-level. In addition, the goal of the chatbot is to provide a correct answer in its first response, without requiring any further interaction.
Correctness differs from the commonly known term hallucination. As hallucinated answers (not grounded in the context) can be correct (Ji et al., 2023a, b). Conversely, non-hallucinated answers can be incorrect if important content is absent (Ji et al., 2023a).
@@
To the best of our knowledge, no prior work defines correctness or not clear enough (Mehri and Eskenazi, 2020; Ji et al., 2023a) for measurement and validation. Therefore, our first objective was to define correctness. To define correctness, we followed a two-pronged approach. First, we looked at 500 chatbot responses for which the support team provided a decision: accept or reject and a short justification for rejection. Further, the first author shadowed (McDonald, 2005) an experienced support employee.
@@
Based on shadowing, discussion, and analysis of rejection reports, the three most common mistakes and, hence, requirements for correctness stood out. They are as defined in the Oxford Dictionary (2024):
@@
"The quality of only saying what is true",
@@
"A close connection with the subject you are discussing or the situation you are in",
@@
"The fact of including all the parts, etc. that are necessary; the fact of being whole".
The annotations often clearly reflected the three requirements. For instance, several comments explicitly indicated issues with completeness, such as: “Incomplete, Jonas (the chatbot) should also provide instructions on how to structure the tasks to proceed earlier.”
A response is considered correct if it contains only true information, as supported by multiple research studies (Li et al., 2018; Wang et al., 2021; Ji et al., 2023b) (truthfulness), is related to the situation and question (relatedness), and comprises all relevant information and solutions (completeness). Later, we designed a plugin soliciting support team responses to understand which requirements the chatbots fall short of, as visualized in Figure 1.
Returning to the literature, we observed that relatedness is well-researched (Zhang et al., 2020; Singh et al., 2021; Yan et al., 2016). However, truthfulness and completeness of generated answers are not. Of the remaining two, we study truthfulness since if an answer is not true, completeness would not matter. For an incomplete answer, the customer can ask follow-up questions, but if untrue information is presented to the customer, it can cause harm to the customer and the company. In the future, a combination of the above three dimensions can be used to measure correctness. The rest of the paper measures truthfulness and assesses it with respect to the manually rated ground truth.
4 Methodology
We combined qualitative and quantitative approaches to define and evaluate answer correctness in the chatbot. First, we analyzed existing feedback, manual annotations, and shadowed a support employee to understand how correctness is assessed in practice (Section 4.1). Based on these insights and literature review, we defined requirements for a correct answer: completeness, relatedness, and truthfulness. We then implemented an extra annotation plugin to collect structured feedback from the support team along these requirements (Section 4.2) and focused on truthfulness for automated evaluation. Using the annotated data, we built a decision tree to model the support team’s assessment process (Section 4.3). This led to the identification of message types (Section 4.4.1), and we derived heuristics describing what makes an answer true (Section 4.4.2). These heuristics guided the selection of literature-based metrics to create automated features (Section 5). The automated features were combined into a truthfulness score (Section 5.2) and validated on a test set (Section 6). Additionally, since most related research and tools are developed for English, we include a side experiment to explore whether using English translations of Dutch responses improves performance.
@@ 4 Methodology
4.1 Defining a right answer
To define what makes an answer right (Section 3), we began by conducting a literature search to identify existing definitions of right chatbot answers. In parallel, we analyzed annotations of the support team of AFAS. These annotations were gathered the months before we started our research. Each response generated by the chatbot could be either accepted or rejected by employees, with a short justification provided for each rejection. In total we were able to collect approximately 500 annotations, providing a rich source of quantitative feedback.
To complement the quantitative analyses, and gain a deeper insight in how support employees assess the rightness of answers, we collected qualitative data through shadowing (McDonald, 2005). The first author shadowed an experienced support employee throughout one workday and interactively discussed the decision-making. The referenced employee has 12 years of experience in various product support areas of the software, has been using chatbot since its launch, and has received relevant AI training, including a course on using ChatGPT. During the discussions, the employee discussed special cases where the chatbot did not meet expectations and what modifications were required to ensure the answer was correct before presenting it to the customer.
The definition of correctness and its requirements were derived by integrating insights from three sources: (1) existing literature and quality metrics, (2) quantitative feedback from 500 chatbot evaluations, and (3) qualitative observations from employee shadowing. We first conducted open coding of the annotations to identify recurrent themes explaining why answers were accepted or rejected. These themes were then compared with concepts identified in prior literature. Through an iterative process of coding, discussion, comparison with existing frameworks, and clustering, we grouped the observed themes into three main requirements: truthfulness, relatedness and completeness. Which together capture the key requirements of a correct chatbot answer. This approach ensured that the resulting definition was both theoretically grounded and empirically validated within the real-world context of customer support.
4.2 Data Gathering
@@
Currently, the developed chatbot of AFAS is utilized by the support team as an assistant. They use the bot to answer questions they have themselves or to answer the questions of a customer. Based on their expert knowledge, they rate the answer of the chatbot for truthfulness. They are encouraged to rate messages, but may choose which messages to rate by themselves. Consequently only a few messages are rated each day, this challenge is posed as we work with a real-case company scenario. The rating is along a Likert scale, which ranges from (1) very untrue to (5) very true. Likert scales are commonly used in similar research by Goodman et al. (2023). While the size of the scale is a topic of debate, a 5-point Likert scale is most commonly used (Borsci et al., 2021; van der Lee et al., 2019; Voigt, 2021). During the research, the team continued to use the chatbot, which meant that new feedback was continually being received.
@@
During the study, the plugin to rate truthfulness is implemented. A few weeks after the rating option is introduced, the data is extracted from the system to form an analysis set, consisting of 79 samples. A few weeks later, a test set is extracted, containing 154 message-answer pairs. For each rated message-answer pair, the context (relevant documents + system prompt) is gathered as well. Finally, to test whether English text performs better than Dutch, we translate the data using Google Translator [^2] https://pypi.org/project/deep-translator/. Since our raters frequently rate an answer as either fully true or fully untrue, the dataset becomes imbalanced, see Table 1. Consequently, we focus less on overall accuracy and more on the accuracy of detecting 1-star and 5-star rated messages. These messages have a larger number and are of greater interest to the company. Since fully untrue answers must not be sent to users, and fully true answers can be sent without human intervention.
@@
Train set 22 2 8 6 41
Test set 32 5 19 16 82
Table 1: Counts of star ratings (1–5), assigned by the support team to reflect message truthfulness, in the training and test datasets.
4.3 Constructing the Tree
@@
To ensure that our metric correlates with human ratings, it is essential to understand how a human determines their rating. In order to represent the thought process of human annotators, we sought a model that could be easily visualized and understood. Furthermore, we sought a method to systematically organize the analysis and maintain a record of the observed characteristics. We opted for a decision tree to depict the mental process, with each node symbolizing the subconscious decisions made by the annotator. We use a manual created tree, as our goal is to mimic the human workflow. If we use automated decision tree builders like Random Forests (Louppe, 2014) or C4.5 (Quinlan, 1993), they would create their own reflection and would not reflect the human workflow.
@@
In relevant literature, automated decision-makers utilize features that range from simple and syntactic to complex and semantic ones. Simple features might verify the presence of keywords (Kumar et al., 2019; Jamil and Hameed, 2023; Roychowdhury et al., 2023), while more complex features involve evaluating the relevance (Merdivan et al., 2020; Zhang et al., 2020) and meaning of a sentence (Vij et al., 2019; Fellbaum and Miller, 1998; Resink, 1995). The decision tree is build up based on this hierarchy, starting with syntactic checks and progressing to nuanced semantic evaluations if the syntactic checks hold. With this approach, the construction and human evaluation process becomes efficient, as it ensures that complex evaluations are not always necessary to be carried out.
@@
The tree is initially constructed by the first author through iterative testing and modification using message-answer pairs. For the initial message-answer pair, the author assessed what makes the answer untrue and added a node describing the mistake. Then, another message-answer pair was evaluated to check if the decision tree properly rejects the answer. If not, a node is added or changed. This procedure is executed for all message-answer pairs and consistently cross-checked with the pairs that have already been evaluated.
While the first version of the decision tree was proposed by the first author, it evolved iteratively to incorporate AFAS-specific context and maintain simplicity. Each of the other two authors contributed this perspective during the iterative process, and the AI development team at AFAS also provided input through discussions on the tree’s construction. Thus, although the tree was primarily built by the first author, it was developed in close consultation with others.
As the decision tree is constructed manually, of course the representation differs if created by another researcher. However, multiple employees of the company confirm that the created decision tree correctly reflects the annotation process. Even when multiple people are involved, the decision tree reflects their interpretation of the mental model, which may differ from the actual mental model. However, by involving several contributors in the process, we believe we reduced this threat and arrived at a more generic and robust mental model. The final version of the decision tree can then be used to determine whether an answer is truthful. Although the decision tree does not make perfect decisions, it effectively visualizes the mental process of an annotator. A part of the tree is shown in Figure 2, and the full tree is included in the replication package (Lassche et al., 2024).
@@
Figure 2: A snippet of the decision tree, indicating whether an answer would be true or not
@@
The first author has no knowledge about the product during the construction of the tree and has relied on the evaluation in the same context as our metric and the LLM. This way, the knowledge of the author is comparable to the knowledge of the LLM and the metric. It is crucial, as the metric, like the author, should be able to calculate a score without a reference answer and, therefore, without relying on external knowledge from such a reference. This is possible since the LLM is not trained, as far as we know on the company documentation. All info in the answer about the product should be contained in the context; otherwise, the LLM will just come up with info on its own. If not in the context, the LLM is assumed to be hallucinating (Roychowdhury et al., 2023).
@@
4.4 Deriving Heuristics from the Decision Tree